OpenStudy (anonymous):

A radioactive material decays at a rate proportional to the amount present. Initially there are 100 milligrams of the material present and after 1 hour(s) the material has 20 milligrams. Find the material decay constant,

OpenStudy (anonymous):

You need to solve:\[\frac{ dm }{ dt }=k·m \rightarrow \frac{ dm }{ m }=k·dt \rightarrow Ln[m(t)]=k·t+A \rightarrow m(t)=e^{A}e^{k·t}=Ce^{k·t}\]\[m(0)=100=C \rightarrow m(t)=100e^{k·t}\]\[m(1)=100e^{k}=20\rightarrow e^k=\frac{ 1 }{ 5 }\rightarrow k=-Ln(5)=-1.609\]Units of "k" are in \[[hours]^{-1}\]

OpenStudy (anonymous):

i didnt mean to close the question

OpenStudy (anonymous):

i had 1.609 but it told me i was wrong

OpenStudy (anonymous):

It has a negative sign.

OpenStudy (anonymous):

i tried both

OpenStudy (anonymous):

then their answer is wrong or you are missing something (do you have to put the units?)

OpenStudy (anonymous):

no i dont need units

OpenStudy (anonymous):

Just try:\[m(t)=100e^{-1.609·t}\]at t=0 and t=1 and check it. Also\[\frac{ dm(t) }{ dt}=-1.609·100·e^{-1.609·t}=-1.609·m(t)\]All conditions are fulfilled with our solution

OpenStudy (anonymous):

Hoow many decimal places you have to put? If it is 2, just put -1.61.If it is one, put -1.6

OpenStudy (anonymous):

is that the mass of the material as a function of time?

OpenStudy (anonymous):

of course. THe problem states that mass decays proportionally to the amount present and that amount has changed in time

OpenStudy (anonymous):

Find the mass of the material N_1 after 5 hours,